OpenStudy (anonymous):

Find the slope and the y intercept of the line. 3x+10y+50

OpenStudy (anonymous):

3 is slope and 10 is y-intercept

OpenStudy (anonymous):

could you rewrite it again, I think you missed the equal sign

OpenStudy (across):

Assuming you meant: \[3x+10y=50\]Then: \[y=-\frac{3}{10}x+5\]

OpenStudy (across):

The slope is: \[-\frac{3}{10}\]The y-intercept is: \[5\]

OpenStudy (anonymous):

3x+10y=50 10y=-3x+50 y=-3/10 x + 5

OpenStudy (anonymous):

I just noticed.. 3x=10y+50

OpenStudy (anonymous):

then 3x=10y+50 3x-50=10y y=3/10 x -5